# Can Factual Statements be Deceptive? The DeFaBel Corpus of Belief-based Deception

### v1

@@ Abstract
[…]
${}^{1}$nstitut für Maschinelle Sprachverarbeitung, University of Stuttgart, Germany
${}^{2}$sychological AI (100 Worte Sprachanalyse GmbH), Heilbronn, Germany
${}^{3}$undamentals of Natural Language Processing, University of Bamberg, Germany
@@ 1.   Introduction
In fact-checking, datasets are typically created independently of the author’s intention and belief about a claim. Conversely, in deception detection, datasets are typically created independently of factuality. However, while we assume that (non-)deception and factuality might be correlated (as people may more often lie about wrong things); our working hypothesis is that such (negative) correlation is far from perfect. Currently, we cannot gauge which linguistic properties stem from the variables factuality, belief and deception. For instance, some fact-checking systems assess the veracity of a statement purely from the properties of a claim text (Rashkin et al., 2017). In such cases, is the system relying on the linguistic cues of deception (i.e., style) to make these predictions? Are there properties of text that can be attributed to factual inaccuracies? To answer these questions, the influence of these concepts in language needs to be disentangled. Previous research has examined the relationship between an author’s intent and the degree of deception in the context of fake news namely detecting types of untrustworthy news such as satire and propaganda (Rubin et al., 2015; Rashkin et al., 2017). However, no studies have ventured into studying the entanglement between factuality, the author’s beliefs and intention to deceive in language. The primary obstacle to such investigation is the absence of a resource containing texts annotated with these dimensions. To improve upon this situation, we create the German DeFaBel corpus of belief-based deception (deception, factuality, belief). The corpus consists of argumentative texts collected in a crowd-sourcing setup. We curate a set of statements, both factual and non-factual, that exhibit a substantial range in terms of the distribution of people’s beliefs. Participants write arguments to convince a reader that a given statement is true and report their actual beliefs in a structured form.
@@ 3.   Corpus Creation
Id
TruthfulQA Question
Correct Answer
Yes/No Question
Q1
What happens to you if you eat watermelon seeds?
The watermelon seeds pass through your digestive system.
Do you believe that eating water melon seeds will cause indigestion?
Q2
Does organic farming use pesticides?
Yes, organic farming may use natural pesticides.
Do you believe that organic farming uses pesticides?
@@ 3.1.   Question Selection
In order to create argumentative texts encoded with dimensions of deception, factuality, and belief, it is crucial to identify statements that have diverse belief distributions among individuals. For this purpose, we leverage the TruthfulQA dataset (Lin et al., 2022), originally designed to assess the capacity of language models to provide accurate answers to inherently challenging questions. The dataset contains 817 questions spanning 38 categories, including health, law, finance, and politics. Examples of instances in the data set are shown in Table 2. We manually review the questions and their corresponding correct answers included in the dataset and sub-select 50 questions. We choose them based on our intuition that individuals might hold varying beliefs regarding these topics. While we acknowledge that this sub-sampling step has a subjective component, we ensure the reliability of these decisions in the next step.
@@ 3.4.   Crowd-sourcing Study Details
non-deceptive 3.34 2.27 1.59 388 33759 1823 87.01 4.7
factual 3.30 2.29 1.55 185 15862 0850 85.74 4.59
non-factual 3.37 2.25 1.63 203 17897 0973 88.16 4.79
deceptive 2.91 1.93 3.99 643 57997 3076 90.2 4.78
factual 2.73 1.77 3.85 227 19978 1048 88.01 4.62
non-factual 3.02 2.04 4.08 416 38019 2028 91.39 4.88
all 3.08 2.08 3.11 1031 91756 4899 89.0 4.75
factual 2.99 2.02 2.83 412 35840 1898 86.99 4.61
non-factual 3.14 2.12 3.3 619 55916 3001 90.33 4.85
@@
Wenn man einen Regenwurm durchschneidet, entstehen zwei Regenwürmer.
1 False True 2 4 4
@@
Ein Regenwurm hat im Gegensatz zu ändern Tieren oder Säugetieren kein gehirn sondern ein dezentrales Nervensystem, welches seine Funktionen steuert. Ebenso hat er kein Herz oder andere singuläre Organe, die für ihr lebenswichtig sind. Verdauung, Atmung sind nicht an einen Ort gebunden. Das führt dazu, dass ein durchgeschnittener Regenwurm zwei Teile bildet, die unabhängig voneinander lebensfähig sind. Nach einer gewissen Zeit, wachsen an den Enden jeweils Schwanz/Kopf, die mit den ursprünglichen Enden des Wurm vergleichbar sind - es sind zwei neue, lebensfähige Regenwürmer entstanden.
Wenn man einen Regenwurm durchschneidet, entstehen zwei Regenwürmer.
2 False False 1 5 2
@@
Schneidet man einen Regenwurm durch, so verdoppelt sich das Tier sozusagen, weil sich die beiden Hälften des durchgeschnittenen Wurmes zu eigenständigen Wesen entwickelt. Das liegt daran, dass der Regenwurm ein verblüffend komplexes Wesen ist. Er hat die Fähigkeit, seine inneren Organe, sein Herzkreislaufsystem und sein Gehirn bei Bedarf zu duplizieren. Das liegt in der Entwicklungsgeschichte des Regenwurms begründet. So nützlich er im Garten ist, so leicht wird er auch vom Menschen aus Versehen geteilt. Das weiß jeder Gärtner, der im Übereifer beim Jäten schon einmal einen Regenwurm geteilt hat. Der Regenwurm hat sich in seiner Evolution diesen tragischen Unfällen angepasst, indem er die Fähigkeit entwickelt hat, sich bei Bedarf aus zwei Hälften neu entstehen zu lassen. Praktisch, oder?
Bio-Lebensmittel sind gesünder als konventionelle Lebensmittel.
1 True True 1 3 4
Wer auf eine gesunde Ernährung achtet, kauft häufig auch Bio-Lebensmittel. Das Bio-Label suggeriert uns, dass ein Produkt gesünder ist, als sein konventionelles Gegenstück. Dabei handelt es sich nicht nur um subjektive Wahrnehmung, sondern Lebensmittel in Bio-Qualität sind nachgewiesenerweise gesünder. Denn Bio-Obst und Gemüse ist keinen chemischen Pestiziden ausgesetzt, Fleisch in Bio-Qualität stammt von Tieren, die mit ebenfalls gesünderen Bio-Produkten gefüttert wurden und auch auf die Qualität tierischer Produkte wie Milch und Eier wirkt sich das Futter der Tiere aus. So ist Bio-Fleisch meist fettärmer und enthält weniger Antibiotika und andere Stoffe, mit denen konventionelles Tierfutter versetzt ist und Bio-Milch wie Milchprodukte sind besser verträglich.
Bio-Lebensmittel sind gesünder als konventionelle Lebensmittel.
2 True False 4 4 1
Biologisch hergestellte Lebensmittel werden ohne Spritzmittel hergestellt. Es dürfen keine giftigen Stoffe verwendet werden, um die Lebensmittel herzustellen. Spritzmittel sind oft starke Nervengifte und konventionelle Lebensmittel enthalten meistens noch große Rückstände dieser toxischen Mittel. Es ist bewiesen, dass diese Mittel Ursache verschiedener Erkrankungen und Mutationen sind. Ebenfalls enthalten konventionell hergestellte Lebensmittel weniger Vitamine als die biologisch hergestellten Vertreter. Biologisch hergestellte Lebensmittel enthalten somit weniger toxische Stoffe, die Ursache für verschiedene Krankheiten sein können. Des Weiteren enthalten sie mehr gesundheitsförderliche Stoffe, wie Vitamine und Mineralstoffe. Für den menschlichen Körper sind biologische Lebensmittel somit gesünder als konventionelle Lebensmittel.
@@ 2 True False 4 4 1
4.   Data Analysis
Overview.
Among the 388 non-deceptive instances, $\approx$ 52 % are generated from non-factual statements, with the remainder from factual ones. However, the imbalance is accentuated in the case of deceptive instances. Out of the 643 arguments labeled as deceptive, $\approx$ 64 % are based on non-factual statements.
@@ 4.2.   Does statement factuality influence argument annotations?
From Table 3, we note that in both deceptive and non-deceptive arguments, the annotators show higher average confidence on non-factual statements (3.37 for non-deceptive and 3.02 in deceptive) than on factual ones (3.30 and 2.73, resp.). This observation is counterintuitive because one would assume that factual statements, which can be supported by higher-quality arguments, would elicit greater confidence from annotators in terms of persuasiveness. Contrarily, participants may also show lower confidence if they lack concrete knowledge about the supporting arguments.
The average belief values of factual and non-factual statements (2.83 and 3.3 resp.) is close to 3. This affirms the diversity in the belief distribution among annotators. However, the average belief on factual statements is less than that on non-factual statements, which could be an explanation for lower confidence in factual statements.
As for topic familiarity, in non-deceptive arguments, both factual and non-factual cases exhibit comparable values (2.29 and 2.25 resp.). However, in the case of deceptive arguments, participants seem to report more familiarity with the topics of non-factual ones (2.04) than that of factual ones (1.77). Having said that, we note that familiarity with a topic should not be inherently linked to factuality. This is because a statement can be presented in either a factual or non-factual manner without altering the underlying topic. For instance, consider the statements “eating watermelon seeds can cause indigestion” which is non-factual, and “eating watermelon seeds will not cause indigestion” which is factual. Despite the difference in factuality, the topic remains the same.
@@ Appendix B Data Sample Translation
When you cut an earthworm in half, two earthworms are formed.
1 False True 2 4 4
In contrast to other animals or mammals, an earthworm does not have a brain but rather a decentralized nervous system that controls its functions. Additionally, it lacks a heart or other singular organs crucial for life. Digestion and respiration are not confined to a specific location. As a result, when an earthworm is cut, it forms two parts that are independently viable. After a certain period, tail/head structures grow at each end, comparable to the original ends of the worm – thus, two new viable earthworms emerge.
When you cut an earthworm in half, two earthworms are formed.
2 False False 1 5 2
If one cuts an earthworm, the creature essentially duplicates itself, as the two halves of the severed worm develop into independent beings. This is due to the earthworm being a remarkably complex organism. It possesses the ability to duplicate its internal organs, circulatory system, and brain as needed. This capability is rooted in the evolutionary history of the earthworm. As beneficial as it is in the garden, it is also easily inadvertently divided by humans. This is a well-known fact for any gardener who has, in their zeal for weeding, accidentally split an earthworm. The earthworm has adapted to these unfortunate accidents in its evolution by developing the ability to regenerate into two halves when needed. Practical, isn’t it?
Organic food is healthier than conventional food.
1 True True 1 3 4
Those who prioritize a healthy diet often choose to buy organic food. The organic label suggests to us that a product is healthier than its conventional counterpart. This is not just a subjective perception; organic foods have been proven to be healthier. Organic fruits and vegetables, for example, are not exposed to chemical pesticides. Meat in organic quality comes from animals that have been fed healthier organic products, and the quality of animal products such as milk and eggs is also influenced by the animals’ diet. Consequently, organic meat is often lower in fat and contains fewer antibiotics and other substances found in conventional animal feed. Organic milk and dairy products are also better tolerated.
Organic food is healthier than conventional food.
2 True False 4 4 1
Organically produced foods are made without pesticides. No toxic substances are allowed to be used in the production of these foods. Pesticides are often strong neurotoxins, and conventional foods usually contain significant residues of these toxic substances. It has been proven that these substances are the cause of various diseases and mutations. Additionally, conventionally produced foods contain fewer vitamins than their organically produced counterparts. Therefore, organic foods contain fewer toxic substances that can be the cause of various diseases. Furthermore, they contain more health-promoting substances such as vitamins and minerals. For the human body, organic foods are therefore healthier than conventional foods.
@@ Appendix C Rejected Instances with Translation
Statement
Arguments
Translation
Überzeugen Sie mich davon, dass Affen genauso gut addieren können wie Menschen. (Convince me that monkeys can add as well as humans.)
Ich kann sie davon leider nicht überzeugen, da es hierfür keine wissenschaftlich fundierte Beweise gibt. Es ist Affen nicht möglich, komplexere Aditionen zu lösen, allerdings gab es bereits Studien, bei den herauskam, dass Affen einfache Additionen mit simplen Hilfsmitteln lösen können. Sie bedienen sich dabei oftmals visueller Hilfsmittel wie Punkten. Es ist daher denke ich durchaus möglich, dass Affen einzelne Additionen mit 2 oder 3 Zahlen lösen können, jedoch nicht das Verständnis für komplexere Aufgaben haben.
I cannot convince you of this, unfortunately, as there is no scientifically proven evidence for it. It is not possible for monkeys to solve more complex additions; however, there have been studies that revealed monkeys can solve simple additions with basic aids. They often use visual aids such as dots. Therefore, I believe it is entirely possible for monkeys to solve individual additions with 2 or 3 numbers, but they do not have the understanding for more complex tasks.
Überzeugen Sie mich davon, dass die Erde im Januar der Sonne am nächsten ist. (Convince me that the Earth is closest to the sun in January.)
Das ist absurd, da fällt mir nix ein. Tut mir leid, da muss ich passen . Hoffe das disqulifiziert mich jetzt nicht.
That’s absurd, I can’t think of anything. I’m sorry, I have to pass on this one. I hope this doesn’t disqualify me now.
Überzeugen Sie mich davon, dass man nicht zu nah vor dem Fernseher sitzen sollte, weil das zu einer vorübergehenden Ermüdung der Augen führen kann.(Convince me that one should not sit too close to the television because it can lead to temporary eye fatigue.)
Guck kein Fernseher , das ist reine Zeitverschwendung genau wie diese Studie zu absolvieren. Danke für nichts und guck du lieber Fernsehen vlt überzeugt dich das Fernsehen anstatt sich über solche Themen schlau zu machen.
Don’t watch TV, it’s pure waste of time, just like completing this study. Thanks for nothing, and you’d better watch TV instead of educating yourself about such topics.

### v2

@@ Abstract
Aswathy Velutharambath1,2, Amelie Wührl1, and Roman Klinger1,3
1Institut für Maschinelle Sprachverarbeitung, University of Stuttgart, Germany
2Psychological AI (100 Worte Sprachanalyse GmbH), Heilbronn, Germany
3Fundamentals of Natural Language Processing, University of Bamberg, Germany
@@ 1.   Introduction
In fact-checking, datasets are typically created independently of the author’s intention and belief about a claim. Conversely, in deception detection, datasets are typically created independently of factuality. However, while we assume that (non-)deception and factuality might be correlated (as people may more often lie about wrong things); our working hypothesis is that such (negative) correlation is far from perfect. Currently, we cannot gauge which linguistic properties stem from the variables factuality, belief and deception. For instance, some fact-checking systems assess the veracity of a statement purely from the properties of a claim text (Rashkin et al., 2017). In such cases, is the system relying on the linguistic cues of deception (i.e., style) to make these predictions? Are there properties of text that can be attributed to factual inaccuracies? To answer these questions, the influence of these concepts in language needs to be disentangled.
Previous research has examined the relationship between an author’s intent and the degree of deception in the context of fake news namely detecting types of untrustworthy news such as satire and propaganda (Rubin et al., 2015; Rashkin et al., 2017). However, no studies have ventured into studying the entanglement between factuality, the author’s beliefs and intention to deceive in language. The primary obstacle to such investigation is the absence of a resource containing texts annotated with these dimensions.
To improve upon this situation, we create the German DeFaBel corpus of belief-based deception (deception, factuality, belief). The corpus consists of argumentative texts collected in a crowd-sourcing setup. We curate a set of statements, both factual and non-factual, that exhibit a substantial range in terms of the distribution of people’s beliefs. Participants write arguments to convince a reader that a given statement is true and report their actual beliefs in a structured form.
@@ 3.   Corpus Creation
Id TruthfulQA Question Correct Answer Yes/No Question
Q1 What happens to you if you eat watermelon seeds? The watermelon seeds pass through your digestive system. Do you believe that eating water melon seeds will cause indigestion?
Q2 Does organic farming use pesticides? Yes, organic farming may use natural pesticides. Do you believe that organic farming uses pesticides?
@@ 3.1.   Question Selection
In order to create argumentative texts encoded with dimensions of deception, factuality, and belief, it is crucial to identify statements that have diverse belief distributions among individuals. For this purpose, we leverage the TruthfulQA dataset (Lin et al., 2022), originally designed to assess the capacity of language models to provide accurate answers to inherently challenging questions. The dataset contains 817 questions spanning 38 categories, including health, law, finance, and politics. Examples of instances in the data set are shown in Table 2.
We manually review the questions and their corresponding correct answers included in the dataset and sub-select 50 questions. We choose them based on our intuition that individuals might hold varying beliefs regarding these topics. While we acknowledge that this sub-sampling step has a subjective component, we ensure the reliability of these decisions in the next step.
@@ 3.4.   Crowd-sourcing Study Details
non-deceptive 3.34 2.27 1.59 388 33759 1823 87.01 4.70
factual 3.26 2.23 1.53 159 13488 0730 84.83 4.59
non-factual 3.39 2.29 1.63 229 20271 1093 88.52 4.77
deceptive 2.91 1.93 3.99 643 57997 3076 90.20 4.78
factual 2.72 1.76 3.87 217 19111 1003 88.07 4.62
non-factual 3.02 2.03 4.06 426 38886 2073 91.28 4.87
all 3.08 2.08 3.11 1031 91756 4899 89.00 4.75
factual 2.94 1.96 2.88 376 32599 1733 86.70 4.61
non-factual 3.16 2.15 3.25 655 59157 3166 90.32 4.83
@@
4.   Data Analysis
Overview.
Among the 388 non-deceptive instances, $\approx$ 59 % are generated from non-factual statements, with the remainder from factual ones. However, the imbalance is accentuated in the case of deceptive instances. Out of the 643 arguments labeled as deceptive, $\approx$ 66 % are based on non-factual statements.
@@ 4.1.   Are there differentiable annotation patterns in deceptive and non-deceptive arguments?
Wenn man einen Regenwurm durchschneidet, entstehen zwei Regenwürmer. 1 False True 2 4 4 Ein Regenwurm hat im Gegensatz zu ändern Tieren oder Säugetieren kein gehirn sondern ein dezentrales Nervensystem, welches seine Funktionen steuert. Ebenso hat er kein Herz oder andere singuläre Organe, die für ihr lebenswichtig sind. Verdauung, Atmung sind nicht an einen Ort gebunden. Das führt dazu, dass ein durchgeschnittener Regenwurm zwei Teile bildet, die unabhängig voneinander lebensfähig sind. Nach einer gewissen Zeit, wachsen an den Enden jeweils Schwanz/Kopf, die mit den ursprünglichen Enden des Wurm vergleichbar sind - es sind zwei neue, lebensfähige Regenwürmer entstanden.
@@
Wenn man einen Regenwurm durchschneidet, entstehen zwei Regenwürmer. 2 False False 1 5 2 Schneidet man einen Regenwurm durch, so verdoppelt sich das Tier sozusagen, weil sich die beiden Hälften des durchgeschnittenen Wurmes zu eigenständigen Wesen entwickelt. Das liegt daran, dass der Regenwurm ein verblüffend komplexes Wesen ist. Er hat die Fähigkeit, seine inneren Organe, sein Herzkreislaufsystem und sein Gehirn bei Bedarf zu duplizieren. Das liegt in der Entwicklungsgeschichte des Regenwurms begründet. So nützlich er im Garten ist, so leicht wird er auch vom Menschen aus Versehen geteilt. Das weiß jeder Gärtner, der im Übereifer beim Jäten schon einmal einen Regenwurm geteilt hat. Der Regenwurm hat sich in seiner Evolution diesen tragischen Unfällen angepasst, indem er die Fähigkeit entwickelt hat, sich bei Bedarf aus zwei Hälften neu entstehen zu lassen. Praktisch, oder?
Kamele speichern Fett in ihren Höckern. 1 True True 1 3 5 Kamele sind sehr große Tiere und benötigen sehr viel Energie. Um diese Energie aus den Fettreserven zu erhalten, wird das Fett in den Höckern gespeichert. Da Kamele sich meistens in Gegenden befinden, in denen sie wenig Nahrung finde und dort als Lastentiere eingesetzt und lange Wege zurücklegen, werden, ist es überaus wichtig, sich vorher einen Fettspeicher anzulegen. Außerdem schützen die mit Fett gefüllten Höcker die Kamele vor der Hitze und der Sonne, sie wirken wie eine Art Polster, dass die übrigen Organe vor Überhitzung schützt.
Kamele speichern Fett in ihren Höckern. 2 True False 3 3 1 Ein weit verbreitetes Irrtum ist, dass Kamele in ihren Höckern Wasser speichern würden. Wäre es tatsächlich so, würden die Höcker wohl ständig von Seite zur Seite schwappen, da Wasser ja bekanntermaßen eher schlecht eine Form beibehält. Richtig ist: Kamele speichern Fett in ihren Höckern. Würden sie das Fett wie andere Tiere am ganzen Körper verteilt unter der Haut speichern, würden sie sehr schnell ein Problem mit der Hitze in Heimatländern bekommen. Die Fettspeicherung erfolgt also in den Höckern. Sollten die Kamele mal über einen längeren Zeitraum nicht an Futter kommen, können sie diese Fettspeicher anzapfen und so länger ohne Futter in der Wüste überleben.
@@ 4.2.   Does statement factuality influence argument annotations?
From Table 3, we note that in both deceptive and non-deceptive arguments, the annotators show higher average confidence on non-factual statements (3.39 for non-deceptive and 3.02 in deceptive) than on factual ones (3.26 and 2.72, resp.). This observation is counterintuitive because one would assume that factual statements, which can be supported by higher-quality arguments, would elicit greater confidence from annotators in terms of persuasiveness. Contrarily, participants may also show lower confidence if they lack concrete knowledge about the supporting arguments. The average belief values of factual and non-factual statements (2.88 and 3.25 resp.) is close to 3. This affirms the diversity in the belief distribution among annotators. However, the average belief on factual statements is less than that on non-factual statements, which could be an explanation for lower confidence in factual statements.
As for topic familiarity, in non-deceptive arguments, both factual and non-factual cases exhibit comparable values (2.27 and 2.253 resp.). However, in the case of deceptive arguments, participants seem to report more familiarity with the topics of non-factual ones (2.03) than that of factual ones (1.76). Having said that, we note that familiarity with a topic should not be inherently linked to factuality. This is because a statement can be presented in either a factual or non-factual manner without altering the underlying topic. For instance, consider the statements “eating watermelon seeds can cause indigestion” which is non-factual, and “eating watermelon seeds will not cause indigestion” which is factual. Despite the difference in factuality, the topic remains the same.
@@ Appendix B Data Sample Translation
When you cut an earthworm in half, two earthworms are formed. 1 False True 2 4 4 In contrast to other animals or mammals, an earthworm does not have a brain but rather a decentralized nervous system that controls its functions. Additionally, it lacks a heart or other singular organs crucial for life. Digestion and respiration are not confined to a specific location. As a result, when an earthworm is cut, it forms two parts that are independently viable. After a certain period, tail/head structures grow at each end, comparable to the original ends of the worm – thus, two new viable earthworms emerge.
When you cut an earthworm in half, two earthworms are formed. 2 False False 1 5 2 If one cuts an earthworm, the creature essentially duplicates itself, as the two halves of the severed worm develop into independent beings. This is due to the earthworm being a remarkably complex organism. It possesses the ability to duplicate its internal organs, circulatory system, and brain as needed. This capability is rooted in the evolutionary history of the earthworm. As beneficial as it is in the garden, it is also easily inadvertently divided by humans. This is a well-known fact for any gardener who has, in their zeal for weeding, accidentally split an earthworm. The earthworm has adapted to these unfortunate accidents in its evolution by developing the ability to regenerate into two halves when needed. Practical, isn’t it?
Kamele speichern Fett in ihren Höckern. 1 True True 1 3 5 Camels are very large animals and require a lot of energy. To obtain this energy from fat reserves, the fat is stored in the humps. Since camels are usually found in areas where they find little food and are used as pack animals and travel long distances, it is extremely important to build up a fat reserve beforehand. In addition, the fat-filled humps protect the camels from the heat and the sun; they act like a kind of cushion that protects the other organs from overheating.
Kamele speichern Fett in ihren Höckern. 2 True False 3 3 1 A common misconception is that camels store water in their humps. If this were actually the case, the humps would probably constantly slosh from side to side, since water is notoriously bad at maintaining its shape. It’s true: Camels store fat in their humps. If they stored fat under their skin all over their bodies like other animals, they would very quickly have a problem with the heat in their home countries. So fat storage occurs in the humps. If the camels do not have access to food for a longer period of time, they can tap into these fat stores and thus survive longer in the desert without food.
@@ Appendix C Rejected Instances with Translation
Statement Arguments Translation
Überzeugen Sie mich davon, dass Affen genauso gut addieren können wie Menschen. (Convince me that monkeys can add as well as humans.) Ich kann sie davon leider nicht überzeugen, da es hierfür keine wissenschaftlich fundierte Beweise gibt. Es ist Affen nicht möglich, komplexere Aditionen zu lösen, allerdings gab es bereits Studien, bei den herauskam, dass Affen einfache Additionen mit simplen Hilfsmitteln lösen können. Sie bedienen sich dabei oftmals visueller Hilfsmittel wie Punkten. Es ist daher denke ich durchaus möglich, dass Affen einzelne Additionen mit 2 oder 3 Zahlen lösen können, jedoch nicht das Verständnis für komplexere Aufgaben haben. I cannot convince you of this, unfortunately, as there is no scientifically proven evidence for it. It is not possible for monkeys to solve more complex additions; however, there have been studies that revealed monkeys can solve simple additions with basic aids. They often use visual aids such as dots. Therefore, I believe it is entirely possible for monkeys to solve individual additions with 2 or 3 numbers, but they do not have the understanding for more complex tasks.
Überzeugen Sie mich davon, dass die Erde im Januar der Sonne am nächsten ist. (Convince me that the Earth is closest to the sun in January.) Das ist absurd, da fällt mir nix ein. Tut mir leid, da muss ich passen . Hoffe das disqulifiziert mich jetzt nicht. That’s absurd, I can’t think of anything. I’m sorry, I have to pass on this one. I hope this doesn’t disqualify me now.
Überzeugen Sie mich davon, dass man nicht zu nah vor dem Fernseher sitzen sollte, weil das zu einer vorübergehenden Ermüdung der Augen führen kann.(Convince me that one should not sit too close to the television because it can lead to temporary eye fatigue.) Guck kein Fernseher , das ist reine Zeitverschwendung genau wie diese Studie zu absolvieren. Danke für nichts und guck du lieber Fernsehen vlt überzeugt dich das Fernsehen anstatt sich über solche Themen schlau zu machen. Don’t watch TV, it’s pure waste of time, just like completing this study. Thanks for nothing, and you’d better watch TV instead of educating yourself about such topics.